OpenStudy (anonymous):

Solve 3(2x) = 7(x−1).

OpenStudy (nincompoop):

what do you think you should do first?

OpenStudy (anonymous):

its x=7

OpenStudy (nincompoop):

you can deal with distributive property first or you can settle the coefficients by making the right hand side and left hand side have the same value of coefficients

OpenStudy (nincompoop):

please do not give out answers

OpenStudy (anonymous):

but if I do the substitution method wouldn't I need to do something else first? @nincompoop

OpenStudy (nincompoop):

substitution method how?

OpenStudy (anonymous):

its more like 3^(2x) = 7^ (x-1)

OpenStudy (nincompoop):

ohhh

OpenStudy (nincompoop):

hahaha

OpenStudy (nincompoop):

did you learn about logarithms yet?

OpenStudy (anonymous):

yes but it makes no sense

OpenStudy (nincompoop):

laughing out loud

OpenStudy (anonymous):

this is serious lol

OpenStudy (anonymous):

I need this class

OpenStudy (anonymous):

Yeah nincompoop.

OpenStudy (nincompoop):

use this \(\large log(a^b) = b~log(a) \)

OpenStudy (nincompoop):